Some European trucks run on energy stored in a rotating flywheel, with an electric motor getting the flywheel up to its top speed of 160 π rad/s. One such flywheel is a solid, uniform cylinder with a mass of 270 kg and a radius of 1.07 m. (a) What is the kinetic energy of the flywheel after charging? (b) If the truck uses an average power of 6.2 kW, for how many minutes can it oQuestion 5 Some European trucks run on energy stored in a rotating flywheel, with an electric motor getting the flywheel up to its top speed of 160 t rad/s. One such flywheel is a solid uniform cylinder with a mass of 270 kg and a radius of 1.07 m. (a) What is the kinetic energy of the flywheel after charging? (b) If the truck uses an average power of 6.2 kW, for how many minutes can it operate between chargings? (a) Number (b) Number Units Unitsperate between chargings?
Answer #1

(a) KE = (1/2)Iw^2 = (1/2)*(1/2)mr^2 w^2

= (1/4)*270*(1.07)^2*(160*pi)^2

= 1.95*10^7 J

(b) t = KE/ P = (1.95*10^7)/(6.2*10^3)

= 3.15*10^3 s